40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13995     Entry time: Thu Jun 21 13:24:00 2018
Author: keerthana 
Type: Update 
Category: elog 
Subject: The cavity scan data of June 20 

(Jon, Keerthana, Sandrine)

We tried to align the AUX and PSL laser yesterday. We collected the data from the spectrum analyser for the Y-ARM reflection and also for the Y-ARM transmission from the ETM mirror. I am attaching the plots here.

Attachment 1: AS110_Beat.pdf  21 kB  Uploaded Thu Jun 21 14:30:00 2018  | Hide | Hide all
AS110_Beat.pdf
Attachment 2: YEND_Beat.pdf  21 kB  Uploaded Thu Jun 21 14:30:11 2018  | Hide | Hide all
YEND_Beat.pdf
ELOG V3.1.3-